Home‎ > ‎Crew 135‎ > ‎

202-GreenHabReport

Officer Poulet to CapCom and to Coordinator Orenstein

The heater has been running at 1500W until 12:30 pm and turned back on
at 5pm.


All plants are looking well.

The pump is operating normally.

 The pepper (?) sprouts mostly look unhealthy. Out of 20, 10 seem not to
have germinated; 5 seem to show a small sprout; 5 look healthy. They
were returned to the hab at 5pm.

Greenhab readings for 5PM today are as follows:

                         Current             Min                 Max
Temperature      48 Degrees F   28 Degrees F  81 Degrees F
Humidity           19%                  14%                 23%

Greenhab readings at 12PM today were as follows:

Meter            0.71 KwH
Duration        19h 16 min

Best regards,

Officer Poulet